OpenStudy (anonymous):

How To Simplify (6i) 3 ?

jhonyy9 (jhonyy9):

3 is exponent ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

do you know how many is i^2 = ? so and i^3 = ?

jhonyy9 (jhonyy9):

and you need calcule again 6^3 =6*6*6 = ?

OpenStudy (anonymous):

I got it now, thank you